OpenStudy (nwolf):

use distributive property 7(w-2)=21

OpenStudy (anonymous):

7w-14=21 21+14=35 7w=35 then just divide which equals 5

OpenStudy (anonymous):

7w-14 = 21 --- distribute the 7 to w-2 7w = 35 ----- add 14 to both sides w = 5 ----- divide both sides by 7

OpenStudy (nwolf):

cool thats what I got w=5